OpenStudy (johnweldon1993):

Having second guesses on this one... Solving the Linear system of Differential Equations

OpenStudy (johnweldon1993):

\[\large x' = \left[\begin{matrix}14 & -16 \\ 12 & -14\end{matrix}\right]x\] Let \(\large a = 14, b = -16, c = 12, d = -14\) \[\large \left\{ x_1' = ax_1 + bx_2\} \\ \{x_2' = cx_1 + dx_2 \right\}\] Derivative of \(x_1'\) \[\large x''_1 = ax_1' + bx'_2\] plugged in \(x'_2 \) from first equation \[\large x''_1 = ax_1' + b(cx_1 + dx_2)\] \[\large x''_1 = ax_1' + bcx_1 + bdx_2\] Now plugged in \( bx_2 \) from first equation \[\large x''_1 = ax_1' + bcx_1 + d(x'_1 - ax_1)\] \[\large x''_1 = ax_1' + bcx_1 + dx'_1 - dax_1\] \[\large x''_1 = (a + d)x_1' + (bc - da)x_1\] Now plugging in numbers \[\large x''_1 = \cancel{(14 -14)x_1'} + ((-16 \times 12) - (-14 \times 14)x_1\] \[\large x''_1 = \cancel{(14 -14)x_1'} + ((-16 \times 12) - (-14 \times 14)x_1\] \[\large x''_1 = (192) - (-196)x_1\] \[\large x''_1 = (-192 + 196)x_1\] \[\large x''_1 = 4x_1\] set = to 0 \[\large x''_1 - 4x_1 = 0\] \[\large = \pm 2 \] so \(\large x_1 = C_1e^{2t} + C_2e^{-2t} \)

OpenStudy (johnweldon1993):

Now take the derivative of that \[\large x_1' = 2C_1e^{2t} - 2C_2e^{-2t}\] plug in \(x'_1\) \[\large 14x_1 - 16x_2 = 2C_1e^{2t} - 2C_2e^{-2t}\] Plug in x_1 we just found \[\large 14(C_1e^{2t} + C_2e^{-2t}) - 16x_2 = 2C_1e^{2t} - 2C_2e^{-2t}\] \[\large 14C_1e^{2t} + 14C_2e^{-2t} - 16x_2 = 2C_1e^{2t} - 2C_2e^{-2t}\] solve for \( x_2 \) \[\large 14C_1e^{2t} + 14C_2e^{-2t} - 2C_1e^{2t} + 2C_2e^{-2t} = 16x_2\] \[\large 12C_1e^{2t} + 16C_2e^{-2t} = 16x_2\] \[\large x_2 = \frac{3}{4}C_1e^{2t} + C_2e^{-2t}\]

OpenStudy (johnweldon1993):

I let \(\large c_1 = 4C_1 \) so \[\large x_1 = 4C_1e^{2t} + C_2e^{-2t}\] \[\large x_2 = 3C_1e^{2t} + C_2e^{-2t}\]

OpenStudy (johnweldon1993):

Typed it out beforehand lol... Does that look about right?

OpenStudy (dumbcow):

yeah everything looks good to me, don't see any mistakes here is what wolfram gives, which is equivalent since constants are arbitrary http://www.wolframalpha.com/input/?i=x%27+%3D+14x+-16y+%2C+y%27+%3D+12x+-14y

OpenStudy (johnweldon1993):

I always forget about wolfram! Thanks @dumbcow

OpenStudy (dumbcow):

haha :)